# MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

### v2

@@ Abstract
Solving mathematical problems using computer-verifiable languages like Lean has significantly impacted mathematical and computer science communities. State-of-the-art methods utilize single Large Language Models (LLMs) as agents or provers to either generate complete proof or perform tree searches. However, single-agent methods inherently lack a structured way to combine high-level reasoning in Natural Language (NL) with Formal Language (FL) verification feedback. To solve these issues, we propose MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought framework, (to the best of our knowledge), the first multi-agent framework for Lean4 theorem proving that balance high-level NL reasoning and FL verification in Long CoT. Using this structured interaction, our approach enables deeper insights and long-term coherence in proof generation, with which past methods struggle. We do this by leveraging emergent formal reasoning ability in Long CoT using our novel LoT-Transfer Learning training-inference pipeline. Extensive experiment shows that our framework achieves 61.07% accuracy rate on the Lean4 version of MiniF2F-Test dataset, largely outperforming GPT-4 (22.95%), single-agent tree search (InternLM-Step-Prover, 50.70%), and whole-proof generation (Godel-Prover, 55.33%) baselines. Furthermore, our findings highlight the potential of combining Long CoT with formal verification for a more insightful generation in a broader perspective.
@@ 1 Introduction
Formal reasoning is a cornerstone of human intelligence and a key objective in machine learning (Newell & Simon, 1956), often evaluated through rigorous mathematical derivations (Yang et al., 2024a). With the rise of Large Language Models (LLMs), Chain-of-Thought (CoT) prompting has emerged as a method to formalize reasoning by generating intermediate steps. This approach not only improves interpretability but also enhances reasoning performance (Wei et al., 2022).
@@
Figure 1: Two main directions of FL theorem proving using LLMs: Single agent tree-search and whole-proof generation with their advantages/disadvantages.
@@
However, the ambiguity of Natural Language (NL) complicates the verification of intermediate steps, particularly in advanced mathematics, where theorem proving is prevalent. This challenge is exacerbated by the growing complexity of modern mathematics, which makes proof verification highly demanding and can lead to errors, as seen in the prolonged validation of Fermat’s Last Theorem (Wang et al., 2024). To address this, researchers propose grounding reasoning in first-order logic, enabling automated verification via Formal Language (FL) verifiers. This framework ensures rigor and has led to the development of tools like Lean (De Moura et al., 2015; Moura & Ullrich, 2021), Isabelle (Paulson, 1994), and HOL Light (Harrison, 2009) for verifiable theorem proving.
@@
Figure 2: MA-LoT Framework: (a) Multi-agent Lean4 theorem proving framework: The LoT-Solver model functions as the prover agent to generate initial Lean4 proofs with emergent NL planning for Lean proof in Long CoT (orange block); then it acts as corrector agent to analyze error from Lean executor in Long CoT to output a refined proof. (b) LoT-Transfer Learning (TL): The novel training-inference pipeline makes formal reasoning ability to emerge in Long CoT (L-CoT) without the need for specifically annotated data. This is achieved by adjusting the system prompt to control the on/off of L-CoT in training and inference. (c) Training procedure of LoT-Solver: We use normal SFT to train general NL L-CoT, use LoT-TL to train on Lean SFT and correction data to make Lean L-CoT emergent capability in LLM.
@@
However, writing mathematical proofs in FL requires significant expertise and effort, as most proofs involve extensive repetition and application of low-resource functions (Jiang et al., 2022). With the rapid progress of LLMs, research has explored LLMs’ application in FL reasoning to automate theorem proving (Polu & Sutskever, 2020; Polu et al., 2022; Jiang et al., 2021, 2022; Yang et al., 2024b; Xin et al., 2024b; Wang et al., 2024; Wu et al., 2024a; Kumarappan et al., 2024; Lin et al., 2024). Prior research follows two main approaches, namely, tree-search (Jiang et al., 2021, 2022; Lin et al., 2024; Xin et al., 2024b; Wang et al., 2024) and whole-proof generation (Polu & Sutskever, 2020; Polu et al., 2022; Yang et al., 2024b; Wu et al., 2024a; Kumarappan et al., 2024). The summary of dis/advantages of these two methods can be found in Figure 1
@@
Tree-search methods train an LLM agent to iteratively generate proof steps by predicting the next tactic based on the current proof state. This is achieved through either direct code writing (Polu & Sutskever, 2020; Polu et al., 2022; Xin et al., 2024b; Wu et al., 2024a; Lin et al., 2024) or retrieval-based techniques (Yang et al., 2024b; Kumarappan et al., 2024). These methods apply FL executor to verify after each step of generation and is able to discover some non-trivial proofs. However, as proof complexity increases, tree-search methods become computationally expensive and lack high-level NL planning to control the overall structure of the proof. This causes the tree-search method unable to find some structured proof that requires high-level analysis of the natural language meaning of the question.
@@
In contrast, whole-proof generation treats theorem proving like the code generation problem, where LLMs generate the entire proof in a single attempt using either supervised training (Wang et al., 2024; Xin et al., 2024b) or prompt engineering (Jiang et al., 2021, 2022). This approach leverages the NL reasoning and high-level planning capabilities of LLMs with predictable computation costs, but it lacks intermediate feedback from FL executors. As a result, the generated proofs often lack post-hoc analysis of errors and tend to perform badly on tedious questions that require non-trivial solutions. In summary, existing single agent (or model) approaches struggle to balance the NL reasoning with the verifiability constraints of FL, motivating the need of a more comprehensive framework.
To address the above challenges, we introduce MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought framework, (to the best of our knowledge), the first multi-agent framework for Lean4 theorem proving. The multi-agent inference method for MA-LoT is shown in Figure 2 (a). MA-LoT framework can find both well-structured and non-trivial proofs by employing a collaborative agent framework and the emergent formal thinking capability in Long CoT. Specifically, the multi-agent framework contains a prover agent to write well-structured proof and a corrector agent to analyze error messages from the Lean executor. The Long Chain-of-Thought (CoT) guides the model to think thoroughly before making outputs. We achieve both agent-type control and high-level NL planning in Long CoT. Additionally, we also improve the self-reflection capability by integrating Lean verification results in Long CoT. The two systems work synergically to achieve higher formal reasoning ability through comprehensive proof planning and systematic error analysis.
To support the MA-LoT framework we designed above, we develop the novel LoT-Transfer Learning (TL) training-inference pipeline to train the LoT-Solver model as shown in Figure 2 (b) & (c). Through this pipeline, formal reasoning ability emerges in Long CoT without needing specifically annotated data. It is done by leveraging transfer learning to integrate capabilities in NL Long CoT reasoning, theorem proof SFT data, and correction data together. The structured adaptation enables the model to be aware of formal states and tactics while maintaining strong NL planning capabilities, leading to more coherent and insightful formal proof generation.
@@
Extensive experiments demonstrate that MA-LoT framework effectively enhances the model’s formal reasoning ability through multi-agent design and emergent formal reasoning in Long CoT. The framework can successfully proof some of the advanced IMO, and AIME problems in the MiniF2F datset (Zheng et al., 2021), with which existing models struggle. Under similar sampling budgets, our framework achieves a 61.07% accuracy rate, surpassing state-of-the-art whole-proof generation models (Godel-Prover (Lin et al., 2025), 55.33%) and tree-search baseline (InternLM-Step-Prover(Wu et al., 2024a), 50.70%).
@@
We summarize our contributions as follows: (1) We introduce MA-LoT (to the best of our knowledge) the first multi-agent framework to balance NL reasoning and FL verification under the Long CoT paradigm for Lean4 theorem proving. (2) We propose the method of using Long CoT to synergically combine the nature of NL and FL, allowing the model to generate in-depth and insightful formal theorem proofs through NL planning and analysis. (3) We develop LoT-TL, a training-inference pipeline that makes field-specific Long CoT capabilities emerge to LLMs without requiring explicitly annotated datasets.
@@
Our framework has broad potential beyond Lean4 theorem proving, demonstrating how formal verification can be effectively integrated with Long CoT reasoning. This approach demonstrates the potential for structured, reflective, and adaptable general text generation through iterative planning and error analysis on formal executors. To accelerate advancements of this field, we plan to open-source our code, dataset, and models in the near future.
@@ 2 Methodology
In this section, we detail the development of the MA-LoT framework and training procedure of LoT-Solver model for Lean4 theorem proving. Our framework is designed to make formal reasoning ability to emerge in Long CoT to achieve deep integration between Natural Language (NL) and Formal Language (FL). This is done under an extreme scarcity of NL-FL aligned data faced by the entire field (Wang et al., 2024). We first outline the preliminaries of LLM formal theorem proving in Section 2.1. Then, we describe the LoT-Transfer Learning (LoT-TL) training pipeline in Section 2.2. Finally, we present comprehensive details on how our trained model facilitates MA-LoT framework for Lean4 proof writing in Section 2.3.
@@ 2.1 Preliminaries
We introduce some preliminary knowledge of applying multi-agent Long Chain-of-Thought (CoT) LLMs to Lean4 formal theorem proving.
Current state-of-the-art methods treat Lean4 code as plain text and input them to LLMs. Some works (Yang et al., 2024b; Wu et al., 2024a) apply LLMs as agents to perform tree-search based on the current proof state. They convert theorem statements and proof states (including premises and goals) into text input for LLMs and ask it to generate the possible next tactic. Other works (Xin et al., 2024b; Wang et al., 2024) treat theorem proving like a code-generation task and develop whole-proof generation methods. This method provides the LLMs with NL instruction, NL theorem statement, and Lean4 formal statement; the intended outcome is the complete Lean4 proof in a single pass, which harnesses the NL reasoning capabilities of LLMs by letting it write the NL plans during proving. The input-output examples for both tree-search and whole-proof generation are presented in Appendix B.
The Long CoT LLMs, represented by O1 (OpenAI, 2024) make long internal NL thinking before outputting the final answer. It largely enhances the NL math reasoning ability of LLMs through self-reflection and correction in Long CoT. However, it still struggles to provide rigorous NL proofs and typically has relatively low FL capability.
Our approach integrates the strengths of both tree-search and whole-proof generation by employing a multi-agent system. We apply Long CoT to regulate the interaction between NL reasoning in LLMs and Lean4 verifier feedback, enabling the model to provide more structured and insightful proofs.
@@ 2.2 LoT-TL Training Pipeline
This section introduces a simple but highly effective training pipeline-LoT-Transfer Learning (TL), which makes Lean4 field-specific Long CoT ability emerge in LLMs without the need for Lean4 Long CoT data. This strategy leverages system prompts to regulate training and inference behaviors; it can be divided into three stages: (1) collecting field-specific Supervised Fine-Tuning (SFT) data (Section 2.2.1), (2) training the model on general natural language Long CoT tasks (Section 2.2.2), and (3) training the model using the transfer learning method on SFT and correction data to make formal Long CoT ability emerge(Section 2.2.3). We use DeepSeek-Prover-v1.5-SFT (Xin et al., 2024b) as the base model. Although we focus on Lean4, our framework shows potential to extend to applying Long CoT reasoning in general fields without requirement for RL or special data annotation.
@@ 2.2.1 Obtain SFT data
The first step of LoT-TL pipeline is to gather a moderate amount of NL-FL aligned SFT data for the specific target field (in our case, Lean4). However, existing open-source datasets do not meet the requirement. They are typically small in size (e.g., MiniF2F (Zheng et al., 2021)), or omit NL annotations (e.g., DeepSeek-Prover-v1 dataset (Xin et al., 2024a)), or exhibit relatively low NL quality (e.g., OBT (Wang et al., 2024)), or lack of Lean4 proofs (e.g., Lean-Workbook (Ying et al., 2024)).
To address this, we compile a new Lean theorem proof dataset named LoT-ProveData (LoT-PD), containing 54,465 data records. Each record contains Lean4 theorem statements, verified proofs with NL explanations as comments, and NL statements. The Lean4 theorem proofs come from two sources: the DeepSeek-Prover-v1 dataset and the annotated Lean-Workbook using TheoremLlama and DeepSeek-Prover-v1.5-RL. Next, inspired by the analysis-then-generate approach in Wang et al. (2023), we employ Qwen-2.5-72B to provide an analysis of Lean4 proofs, followed by writing NL proof based on the analysis. Finally, we integrate these NL proofs as comments in the Lean4 code by Qwen. For data records lacking NL statements, we generate NL statements using a similar method. The core components of our LoT-ProveData is:
@@
During proof generation for the ProverData, some incorrect proofs were also produced. We recorded these alongside their error messages to form LoT-CorrectionData (LoT-CD), consisting of 64,912 records of correct-incorrect Lean4 proofs plus associated error messages and NL statement and proof. LoT-CD is used to train the model’s error analysis and correction capability. The core part of the LoT-CorrectionData
{FL Statement, Correct FL proof, Error messages, Incorrect FL proof, NL statement} These datasets work together to improve both the prover and corrector agent’s capability, acting as the source of strong NL-FL joint thinking ability for our multi-agent framework.

### v3

# MA-LoT: Model-Collaboration Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

@@ Abstract
Solving mathematical problems using computer-verifiable languages like Lean has significantly impacted the mathematical and computer science communities. State-of-the-art methods utilize a single Large Language Model (LLM) to generate complete proof or perform tree search, but they fail to balance these tasks. We propose MA-LoT: Model-CollAboration Lean-based Long Chain-of-Thought, a comprehensive framework for Lean4 theorem proving to solve this issue. It separates the cognition tasks of general NL for whole-proof generation and error analysis for proof correction using the model-collaboration method. We achieve this by structured interaction of the LLM and Lean4 verifier in Long CoT. To implement the framework, we propose the novel LoT-Transfer Learning training-inference pipeline, which enables the Long CoT thinking capability to LLMs without special data annotation. Extensive experiment shows that our framework achieves a 61.07% accuracy rate on the Lean4 version of the MiniF2F-Test dataset, largely outperforming DeepSeek-V3 (33.61%), single-model tree search (InternLM-Step-Prover, 50.70%), and whole-proof generation (Godel-Prover, 55.33%) baselines. Furthermore, our findings highlight the potential of combining Long CoT with formal verification for a more insightful generation in a broader perspective.
@@ 1 Introduction
Formal reasoning is a cornerstone of human intelligence and a key objective in machine learning (Newell & Simon, 1956), often evaluated through rigorous mathematical derivations (Yang et al., 2024a). With the rise of Large Language Models (LLMs), Chain-of-Thought (CoT) prompting has emerged to formalize reasoning by generating intermediate steps. This approach improves interpretability and enhances reasoning performance (Wei et al., 2022).
@@
Figure 1: Two main directions of FL theorem proving using LLMs: Single model tree-search and whole-proof generation with their advantages/disadvantages.
@@
However, the ambiguity of Natural Language (NL) complicates verifying intermediate steps, particularly in advanced mathematics, where there is no answer to check but theorems to prove. This challenge is exacerbated by the growing complexity of modern mathematics, which makes proof verification highly demanding and can lead to errors, as seen in the prolonged validation of Fermat’s Last Theorem (Wang et al., 2024). Researchers propose grounding reasoning in rigorous first-order logic to address this, enabling automated verification via Formal Language (FL) verifiers. This framework ensures rigor and has led to the development of tools like Lean (De Moura et al., 2015; Moura & Ullrich, 2021), Isabelle (Paulson, 1994), and HOL Light (Harrison, 2009) for verifiable theorem proving.
@@
Figure 2: MA-LoT Framework: (a) Model-collaboration Lean4 theorem proving framework: The LoT-Solver model functions as the prover to generate initial Lean4 proofs with emergent NL planning for Lean proof in Long CoT (orange block); then it acts as corrector to analyze error from Lean executor in Long CoT to output a refined proof. (b) LoT-Transfer Learning (TL): The novel training-inference pipeline enables formal reasoning ability to emerge in Long CoT (L-CoT) without the need for specifically annotated data. This is achieved by adjusting the system prompt to control the on/off of L-CoT in training and inference. (c) Training procedure of LoT-Solver: We use normal SFT to train general NL L-CoT, use LoT-TL to train on the Lean SFT, and correction data to make Lean L-CoT an emergent capability in LLM.
@@
However, writing mathematical proofs in FL requires significant expertise and effort, as most proofs involve extensive repetition and application of low-resource functions (Jiang et al., 2022). With the rapid progress of LLMs, research has explored LLMs’ application in FL reasoning to automate theorem proving (Polu & Sutskever, 2020; Polu et al., 2022; Jiang et al., 2021, 2022; Yang et al., 2024b; Xin et al., 2024b; Wang et al., 2024; Wu et al., 2024a; Kumarappan et al., 2024; Lin et al., 2024). Prior research follows two main approaches, namely, tree-search (Jiang et al., 2021, 2022; Lin et al., 2024; Xin et al., 2024b; Wang et al., 2024) and whole-proof generation (Polu & Sutskever, 2020; Polu et al., 2022; Yang et al., 2024b; Wu et al., 2024a; Kumarappan et al., 2024). The summary of dis/advantages of these two methods can be found in Figure 1.
@@
Tree-search methods train an LLM to iteratively generate proof steps by predicting the next tactic based on the current proof state. This is achieved through either direct code writing (Polu & Sutskever, 2020; Polu et al., 2022; Xin et al., 2024b; Wu et al., 2024a; Lin et al., 2024) or retrieval-based techniques (Yang et al., 2024b; Kumarappan et al., 2024). Tree-search methods apply an FL executor to verify after each generation step and can discover some non-trivial proofs. However, as proof complexity increases, tree-search methods become computationally expensive and lack high-level NL planning to control the overall structure of the proof. The lack of overall planning causes the tree-search method to be unable to find some structured proof that requires high-level NL planning to answer the question.
@@
In contrast, whole-proof generation treats theorem proving like the code generation problem, where LLMs generate the entire proof in a single attempt using either supervised training (Wang et al., 2024; Xin et al., 2024b) or prompt engineering (Jiang et al., 2021, 2022). This approach leverages LLMs’ NL reasoning and high-level planning capabilities with predictable computation costs, but it lacks intermediate feedback from FL executors. As a result, the generated proofs often lack post-hoc analysis of errors and tend to perform badly on tedious questions that require non-trivial or tedious tricks. In summary, existing single-model approaches struggle to balance the NL reasoning with the FL verifiability, motivating the need for a more comprehensive framework.
To address the above challenges in Lean4 theorem proving, we introduce MA-LoT: Model-CollAboration Lean-based Long Chain-of-Thought, a comprehensive framework designed to coordinate multiple models for effective formal reasoning. As illustrated in Figure 2 (a), MA-LoT separates the cognitive process into two distinct tasks: high-level proof planning and writing, as well as fine-grained correction based on verifier feedback. This separation is enabled by a collaborative model architecture and the emergent reasoning capability of Long CoT, which was developed during training. The framework consists of two core components: a prover model, responsible for generating well-structured proofs, and a corrector model, which interprets feedback from the Lean4 verifier to refine the output. Both models are enhanced with Lean4 field-specific Long CoT capability, allowing them to reason more thoroughly before producing outputs. Moreover, we integrate Lean verification results into Long CoT to improve the system’s self-reflection and correction abilities.
To implement the MA-LoT framework, we propose a novel training-inference pipeline called LoT-Transfer Learning (LoT-TL), which is used to train the LoT-Solver model, as shown in Figure 2 (b) and (c). This pipeline enables the emergence of using Long CoT to solve the FL problems without requiring specifically annotated formal data. It is achieved by leveraging transfer learning across three data sources: NL-based general Long CoT reasoning, SFT data for theorem proving under the code completion task, and in/correct proof pairs based on Lean4 feedback. Through structured adaptation, LoT-TL equips the model with awareness of FL proof states and tactics while preserving its strong NL reasoning and planning capabilities. Then, the model trained under this pipeline can generate coherent and insightful formal proofs grounded in both symbolic precision and high-level planning.
@@
Extensive experiments demonstrate that MA-LoT framework effectively enhances the model’s formal reasoning ability through model-collaboration design and emergent formal reasoning in Long CoT. The framework can successfully prove some of the advanced IMO and AIME problems in the MiniF2F dataset (Zheng et al., 2021), with which existing models struggle. Under similar sampling budgets, our framework achieves a 61.07% accuracy rate, surpassing state-of-the-art whole-proof generation models (Godel-Prover (Lin et al., 2025), 55.33%) and tree-search baseline (InternLM-Step-Prover(Wu et al., 2024a), 50.70%).
@@
We summarize our contributions as follows: (1) We introduce MA-LoT, a comprehensive model-collaboration framework to balance NL reasoning and FL verification under the Long CoT paradigm for Lean4 theorem proving. (2) We propose using Long CoT to synergically combine the nature of NL and FL, allowing the model to generate in-depth and insightful formal reasoning through NL planning and analysis. (3) We develop LoT-TL, a training-inference pipeline that makes field-specific Long CoT capabilities emerge to LLMs without requiring explicitly annotated datasets.
@@
Our framework has broad potential beyond Lean4 theorem proving, demonstrating how formal verification can be effectively integrated with Long CoT reasoning. This approach reveals the potential for structured, reflective, and adaptable general text generation through iterative planning and error analysis on formal executors. To accelerate advancements in this field, we open-source our code at https://github.com/RickySkywalker/LeanOfThought-Official
@@ 2 Methodology
In this section, we detail the development of the MA-LoT framework and training procedure of LoT-Solver model for Lean4 theorem proving. Our framework’s core idea is to enable the model to perform Long Chain-of-Thought (CoT) reasoning under the context of Lean4 theorem proving. The Long CoT is designed to perform deep integration between Natural Language (NL) and Formal Language (FL) reasoning. Training of such Long CoT thinking model is achieved under an extreme scarcity of NL-FL aligned data faced by the entire field (Wang et al., 2024). We introduce the methods by first outlining the preliminaries of LLM formal theorem proving in Section 2.1. Then, we describe the LoT-Transfer Learning (LoT-TL) training pipeline in Section 2.2. Finally, we present comprehensive details on how our trained model facilitates MA-LoT framework for Lean4 proof writing in Section 2.3.
@@ 2.1 Preliminaries
We introduce preliminary knowledge of applying Long Chain-of-Thought (CoT) LLMs to Lean4 formal theorem proving in a model-collaboration manner.
Current state-of-the-art methods treat Lean4 code as plain text and feed it directly to LLMs. There are two main branches of techniques to apply LLMs for Lean4 theorem proving. The first branch is the tree-search method (Yang et al., 2024b; Wu et al., 2024a). This method converts the Lean4 theorem statement and current proof state (including conditions and goals based on the existing proofs) into plain text as input to LLMs and asks it to generate the next possible tactic to complete the proof. The other direction is whole-proof generation methods. It provides the LLMs with NL instruction, NL theorem statement, and Lean4 statement to the LLMs. The intended outcome is to generate a complete Lean4 proof in a single pass. This is achieved by first leveraging the LLM’s NL reasoning to produce a high-level plan, then guiding the generation of the actual Lean4 proof The input-output examples for both tree-search and whole-proof generation are presented in Appendix B.
The Long CoT LLMs, represented by O1 (OpenAI, 2024), make long internal NL thinking before outputting the final answer. It largely enhances the NL math reasoning ability of LLMs through self-reflection and correction in Long CoT. However, it still struggles to provide rigorous NL proofs and typically has relatively low FL capability.
Our approach combines the strengths of tree search and whole-proof generation methods through a model-collaboration system. By leveraging Long CoT, we coordinate the interaction between NL and FL in LLMs, which allows the model to write more structured and insightful proofs.
@@ 2.2 LoT-TL Training Pipeline
This section introduces a simple but effective training pipeline LoT-Transfer Learning (LoT-TL). The pipeline intends to make LLMs have the ability to perform Lean4 field-specific Long CoT reasoning. Such ability can be obtained through LoT-TL without the need for specifically annotated Long CoT data. The main idea of LoT-TL is to leverage system prompts to regulate training and inference behaviors, which can be divided into the following stages: (1) collecting field-specific Supervised Fine-Tuning (SFT) data (Section 2.2.1), (2) training the model on general natural language Long CoT tasks (Section 2.2.2), and (3) training the model using the transfer learning method on SFT and correction data to make formal Long CoT ability emerge(Section 2.2.3). Although this paper focuses on Lean4, our framework shows potential to be applied to more fields, making LLMs obtain field-specific Long CoT capability without RL or special data annotation.
@@ 2.2.1 Obtain SFT data
The first step of LoT-TL pipeline is to gather a moderate amount of NL-FL aligned SFT data for the specific target field (in our case, Lean4). However, existing open-source datasets do not meet the requirement. They are typically small in size (e.g., MiniF2F (Zheng et al., 2021)), or omit NL annotations (e.g., DeepSeek-Prover-v1 dataset (Xin et al., 2024a)), or exhibit relatively low NL quality (e.g., OBT (Wang et al., 2024)), or lack Lean4 proofs (e.g., Lean-Workbook (Ying et al., 2024)).
To address this, we compile a new Lean theorem proof dataset named LoT-ProveData (LoT-PD), containing 54,465 data records. Each record contains Lean4 theorem statements, verified proofs with NL explanations as comments, and NL statements. The Lean4 theorem proofs come from two sources: the DeepSeek-Prover-v1 dataset and the annotated Lean-Workbook using TheoremLlama and DeepSeek-Prover-v1.5-RL. Next, inspired by the analysis-then-generate approach in Wang et al. (2023), we employ Qwen-2.5-72B to provide an analysis of Lean4 proofs, followed by writing the NL proof based on the analysis. Finally, we integrate these NL proofs as comments in the Lean4 code by Qwen. For data records lacking NL statements, we generate NL statements using a similar method. The core components of our LoT-ProveData are as follows:
@@
During proof generation for the ProverData, some incorrect proofs were also produced. We recorded these alongside their error messages to form LoT-CorrectionData (LoT-CD), consisting of 64,912 records of correct-incorrect Lean4 proof pairs with error messages of the incorrect proof. Additionally, it contains an NL statement and proof of the theorem. LoT-CD is used to train the model’s error analysis and correction capability. The core part of the LoT-CD is:
{FL Statement, Correct FL proof, Error messages, Incorrect FL proof, NL statement} These datasets work together to improve both the prover and corrector models’ capability, acting as the source of strong NL-FL joint thinking ability for our model-collaboration framework.